1)If the sum of the areas of two circles with radii r1 and r2 is equal to the area of
a circle of radius r, then
a) r1 +r2 = r b) r1^2 +r2^2 = r^2 c) r1 +r2 < r d) r1^2 +r2^2< r^2
2) 2) If the sum of the circumferences of two circles with radii R1 and R2 is equal to the
circumference of a circle of radius R, then

(A) R1 + R2 = R (B) R1 + R2 > R (C) R1 + R2 < R
(D) Nothing definite can be said about the relation
among R1, R2 and R

Answer by edjones(8007)

Let pi= the constant c
Let r1=rx, r2=ry and r=rz (less confusion for me)
1)
c*r^2=A
Ax+Ay=Az
cx^2+cy^2=cz^2
b) x^2+y^2=z^2 divide each side by c
.
2)
2cr=C
2cx+2cy = 2cz
a) x+y=z divide each side by 2c.
.
Ed